OpenStudy (anonymous):

x^8=y^2+87 Find all the values of x and y for which x and y are prime numbers.

OpenStudy (anonymous):

..........

OpenStudy (anonymous):

hehehehe u thought u where gonna get the answer xD

OpenStudy (anonymous):

Sorry, but I know the answer as this question is mine.

OpenStudy (anonymous):

Oh... What an easy question..! I don't want to solve it.. Ha ha ha ha...

OpenStudy (anonymous):

ahahahaha

OpenStudy (anonymous):

Sorry the question is little changed.

OpenStudy (anonymous):

is this allowed ? (x^4-y)(x^4+y) = 87 since it's either 87*1 or 29*3 we need x^4-y =1 and x^4+y=87 we wont get here prime number for x or x^4-y=3 x^4+y=29 2x^4 = 32 x^4 = 16 x = 2 y = 13 so only x = 2 and y = 13 ?

OpenStudy (anonymous):

Ya...... that looks correct.